OpenStudy (anonymous):

An exterior angle of a regular polygon measures 30 degrees. What is the measure of an interior angle?

jimthompson5910 (jim_thompson5910):

Measure of Interior Angle = 180 - Measure of Exterior Angle Measure of Interior Angle = 180 - 30 Measure of Interior Angle = 150 degrees

OpenStudy (campbell_st):

the sum of the exterior angles of a polygon equal 360 degrees so if an exterior angle is 30 then the number of sides or angles in the polygon is found by 360/30 = 12 the polygon has 12 sides next the angle sum of a polygon is (n - 2)*180 n = the number of sides substitute 12 into the formula to find the angle sum once you know that... divide the angle sum by 12 to get the size of 1 interior angle.